The weight of your 1420 kg car is supported equally by its four tires, each inflated to a gauge pressure of 32.0 lb/in2 .

What is the area of contact each tire makes with the road?

If the gauge pressure is increased, does the area of contact increase, decrease, or stay the same?

What gauge pressure is required to give an area of contact of 124 cm2 for each tire?
Answer #1

Pressure=Force/Area

1 DATE PAGE NO Sol. Mass of car, M = 1420kg | Weight W = Mg = 1420 X9.8 = 13916N Pressure on each tyse AP = f - A p = W/4 lu